OpenStudy (anonymous):

h(x) = 1/2 cos (-Πx)

OpenStudy (anonymous):

Is the amplitude 0.5 and the period 2 ?

zepdrix (zepdrix):

\[\Large\rm period=\frac{2\pi}{\pi}\]Mmmm yes, looks good! :)